OpenStudy (anonymous):

How can I simplify this.. 45^(1/2) -5

OpenStudy (anonymous):

remember \[x^{1/2}=\sqrt{x}\] so \[45^{1/2}=\sqrt{45}\]

OpenStudy (anonymous):

Yeah I know that. But how to solve..

OpenStudy (anonymous):

Are there any easy square numbers that are factors of 45?

OpenStudy (anonymous):

\[\sqrt{45}\] = \[\sqrt{9\times5} = \sqrt{3^{2}\times5} = 3\sqrt{5}\] what do you mean like this?

OpenStudy (anonymous):

\[3\sqrt{5}-5= what\]

OpenStudy (anonymous):

That's all.. you are just simplifying

OpenStudy (anonymous):

Can't we simplify further

OpenStudy (anonymous):

It is similar to simplifying fractions... at the end, you just have a more simple expression than you started with

OpenStudy (anonymous):

Yeah I agree with u all

OpenStudy (anonymous):

how about this \[3\sqrt{5}-5=x\]

OpenStudy (anonymous):

What do you mean? Are you asking whether you can express your response like that? Or are you asking whether you can simplify further?

OpenStudy (anonymous):

Can\[(3\sqrt{5}-5)^{2}=x ^{2}\]'t we remove the root by squaring both sides

OpenStudy (anonymous):

by squaring both sides, you have two terms inside the parenthesis. You can't just apply the "square" to the first term to remove the root, because that expression is really:\[(3\sqrt{5} - 5)(3\sqrt{5}-5) = (3\sqrt{5})^{2}-2(5)(3\sqrt{5)}) +\]

OpenStudy (anonymous):

oops, left off the 25 off the end of that

OpenStudy (anonymous):

It all simplifies down the same anyway, except with other unintended problems that don't matter for this one. Your earlier answer may not look "done", but it was really correct... completely simplified

OpenStudy (anonymous):

Ohh thank U very much for your respond

OpenStudy (anonymous):

Squaring both sides for something like this is probably a bad instinct... @completeidiot might be able to tell you why

OpenStudy (anonymous):

ug what a mess this has become we started with this \[45^{1/2} -5= \sqrt{45}-5\] all you really need to do is simplify the square root getting \[3\sqrt5-5\] this is your simplified expression it doesnt need to be equal to anything remember that an expression is not the same as an equation

OpenStudy (anonymous):

an expression is very different from an equation if you're asked to simplify an expression, the result will still be an expression remember that an equation has an equal sign in it why squaring both sides is a bad idea with that equation simple answer, you still get sqrt(5) on one side which doesnt really help and if x=35√−5 why would you need to do anything further if you already know x

OpenStudy (anonymous):

long story short expressions dont need to be equal to anything

OpenStudy (anonymous):

Yeah I agree with u I didn't know abt the difference between an expression an an equation thnx for explaining abt this Thnx very much I really appreciate both of u @completeidiot & @JakeV8 V8

OpenStudy (anonymous):

\[\sqrt{45}=3\sqrt{5}=6.7082 is simplfied answer i think this on minus with f gives u 1.7082

OpenStudy (anonymous):

Unless they are asking you to run it through a calculator, typically you can leave it in a simplified form that includes a root. If you need the decimals, by all means, use the calculator, but it's just two different forms.